### v3

@@ 8.2 Simulation study
To simulate ordinal variables that were generated by a nonnormal latent distribution $G$, we employ the VITA simulation method of Grønneberg & Foldnes, (2017). For a pre-specified value of the population correlation $\rho_{G}$, the VITA method models the latent random vector $(\xi,\eta)$ such that the individual variables $\xi$ and $\eta$ both possess standard normal marginal distributions with population correlation set equal to $\rho_{G}=\mathrm{Cor}_{G}\left[\xi,\ \eta\right]$, but are not jointly normally distributed. Instead, their joint distribution $G$ is equal to a pre-specified nonnormal copula distribution, such as the Clayton or Gumbel copula. Grønneberg & Foldnes, (2017) show that discretizing such VITA latent variables yields ordinal observations that could have not been generated by a standard bivariate normal distribution, thereby ensuring proper violation of the latent normality assumption.
@@
Figure 8: Boxplot visualization of the bias of the robust and ML estimator for the polychoric correlation coefficient, $\widehat{\rho}_{N}-\rho_{G}$, under distributional misspecification through a Clayton copula with correlation $\rho_{G}=0.9$ (left) and $\rho_{G}=0.3$ (right), across 1,000 repetitions. The tuning constant of the robust estimator is set to $c=0.6$.
@@
such that both resulting ordinal variables have five response options each. We generate $N=1,000$ ordinal responses according to this data generation process and compute across 1,000 repetitions the same estimators and performance measures as in the simulations in Section 6.
@@
Figure 8 provides a boxplot visualization of the bias of each estimator under both Clayton copulas across the repetitions. [^7] In seven of the 1,000 repetitions for the Clayton copula with correlation 0.9, the robust estimator experienced numerical instability and subsequently did not converge to a solution. Consequently, these unconverged estimates are omitted from Figure 8. We explain this numerical issue in detail in Appendix B. For correlation 0.9, the ML estimator exhibits a noteworthy bias, whereas the robust estimator remains accurate, albeit with a larger estimation variance as compared to partial misspecification (cf., Section 6). Conversely, for the weaker correlation 0.3, both estimators are fairly accurate with median biases of about 0.025.
@@
These results suggest that the Clayton copula with correlation 0.9 might be reasonably approximable by a mixture of a normal distribution that covers the center of the probability mass and some nonnormal distribution that covers the tails (cf., Section 8.1). Indeed, Figure 7 indicates that the normal distribution and Clayton copula at correlation 0.9 seem to behave similarly in the center, but deviate from one another towards the tails. On the other hand, at correlation 0.3, the two densities do not appear to be drastically different from one another, which may explain why both ML and robust estimator work reasonably well for such distributional misspecification: If the latent distribution is not too different from a normal distribution, then the polychoric model may offer a satisfactory fit despite being technically misspecified.
@@
Overall, this simulation study demonstrates that in some cases of distributional misspecification, robustness can be gained with our estimator, compared to ML estimation. However, it also demonstrates that there are cases of distributional misspecification for which the ML estimator still works quite well such that the robust estimator offers little gain. Nevertheless, the fact that in some situations robustness can be gained under distributional misspecification represents an overall gain in robustness.
@@ 9 Discussion and conclusion
We consider a situation where the polychoric correlation model is potentially misspecified for a subset of responses in a sample, that is, a set of uninformative observations not generated by a latent standard normal distribution. This model misspecification framework, called partial misspecification here, stems from the robust statistics literature, and a relevant special case is that of careless respondents in questionnaire studies. We demonstrate that maximum likelihood estimation is highly susceptible to the presence of such uninformative responses, resulting in possibly large estimation biases. As a remedy, we propose an estimator based on Welz, (2024) that is designed to be robust to partial model misspecification. Our estimator generalizes maximum likelihood estimation, does not make any assumption on the magnitude or type of potential misspecification, comes at no additional computational cost, and possesses attractive statistical guarantees such as asymptotic normality. It furthermore allows to pinpoint the sources of potential model misspecification through the notion of Pearson residuals (PRs). Each possible response option is assigned a PR, where values substantially larger than the ideal value 0 imply that the response in question cannot be fitted well by the polychoric correlation model. In addition, the methodology proposed in this paper is implemented in the free open source package robcat for the statistical programming environment R and is publicly available at https://github.com/mwelz/robcat.
@@
Although not covered by our partial misspecification framework, we also discuss how and when our estimator can offer a robustness gain (compared to ordinary maximum likelihood) when the polychoric model is misspecified for all observations in a sample, which has been a subject of interest in recent literature. In essence, there can be a robustness gain if the latent nonnormal distribution that generated the data can be reasonably well approximated by a mixture of a normal distribution that covers the center of probability mass and some unspecified nonnormal distribution that covers the tails.
@@
We verify the enhanced robustness and theoretical properties of our robust generalized maximum likelihood estimator in simulation studies. Furthermore, we demonstrate the estimator’s practical usefulness in an empirical application on a Big Five administration, where we find compelling evidence for the presence of careless respondents as source of partial model misspecification.
@@
However, our estimator depends on a user-specified choice of a tuning constant $c$, which governs a tradeoff between robustness and efficiency in case of misspecification. While simulation experiments suggest that the choice $c=0.6$ provides a good tradeoff and estimates do not change considerably for other finite choices of $c$, a detailed investigation on this tuning constant needs to be carried out in future work. As an alternative, one could consider other discrepancy functions that do not depend on tuning constants, like the ones discussed in Welz, (2024).
@@
The methodology proposed in this paper suggests a number of extensions. For instance, one could use a robustly estimated polychoric correlation matrix in structural equation models to robustify such models against misspecification. Similar robustification could by achieved in, for instance but not limited to, principal component analyses, multidimensional scaling, or clustering. In addition, Welz, (2024) proposes a general statistical test to pinpoint possible sources of model misspecification on the individual response level. This test could be used to test in a statistically sound way whether a given response can be fitted well by the polychoric correlation model. We leave these avenues to further research.
@@ Appendix A Asymptotic properties of the robust estimator
This appendix section presents the limit theory of the proposed robust estimator $\widehat{\bm{\theta}}_{N}$ in (10). Throughout this section, we assume that the number of response categories, $K_{X}$ and $K_{Y}$, are fixed and known, and that the sample $\{(X_{i},Y_{i})\}_{i=1}^{N}$ used to compute an estimate $\widehat{\bm{\theta}}_{N}$ has been generated by the process in (1) where the latent $\{(\xi_{i},\eta_{i})\}_{i=1}^{N}$ are draws from distribution $G_{\varepsilon}$ in (7) with unobserved misspecification fraction $\varepsilon\in[0,0.5)$ and unknown misspecifying distribution $H$.
@@ A.1 Main theorem
We start by introducing additional notation.
Define the $d$-dimensional gradient vector of $\log\left(p_{xy}(\bm{\theta})\right)$ for response $(x,y)\in\mathcal{X}\times\mathcal{Y}$ at parameter $\bm{\theta}\in\bm{\Theta}$ as
$\bm{s}_{xy}(\bm{\theta})=\frac{1}{p_{xy}(\bm{\theta})}\left(\frac{\partial}{\partial\bm{\theta}}\ p_{xy}(\bm{\theta})\right).$
Further, given tuning constant $c\in[0,\infty]$, define the $d\times d$ matrix
$\bm{A}_{xy}(\bm{\theta})=\Bigg{(}\frac{f_{\varepsilon}\left(x,y\right)}{p_{xy}(\bm{\theta})}\mathds{1}\left\{\frac{f_{\varepsilon}\left(x,y\right)}{p_{xy}(\bm{\theta})}-1\in[-1,c]\right\}+(c+1)\mathds{1}\left\{\frac{f_{\varepsilon}\left(x,y\right)}{p_{xy}(\bm{\theta})}-1\ >c\right\}\Bigg{)}\frac{\partial^{2}}{\partial{\bm{\theta}}\partial{\bm{\theta}}^{\top}}p_{xy}(\bm{\theta})$
@@
For $c\in[0,\infty]$ the tuning constant in the discrepancy function, assume that $\frac{f_{\varepsilon}\left(x,y\right)}{p_{xy}(\bm{\theta}_{0})}-1\neq c$ for any $(x,y)\in\mathcal{X}\times\mathcal{Y}$. Then, under certain regularity conditions that do not restrict the degree or type of possible misspecification of the polychoric model beyond $\varepsilon\in[0,0.5)$, when $N\to\infty$ it holds true that
@@
$\bm{U}\left(\bm{\theta}\right)$ $=\bm{W}\left(\bm{\theta}\right)\bm{\Omega}\bm{W}\left(\bm{\theta}\right)^{\top}\qquad\text{and}$
$\bm{M}\left(\bm{\theta}\right)$ $=\sum_{x\in\mathcal{X}}\sum_{y\in\mathcal{Y}}\Bigg{[}\mathds{1}\left\{\frac{f_{\varepsilon}\left(x,y\right)}{p_{xy}(\bm{\theta})}-1\in[-1,c]\right\}\bm{s}_{xy}(\bm{\theta})\bm{s}_{xy}(\bm{\theta})^{\top}-\bm{A}_{xy}(\bm{\theta})\Bigg{]}$
@@
It can be shown that in the absence of misspecification (such that $\bm{\theta}_{0}=\bm{\theta}_{*}$), the asymptotic covariance matrix $\bm{\Sigma}\left({\bm{\theta}_{*}}\right)$ is equal to the inverted Fisher information matrix of the polychoric model (Lemma B.5 in Welz,, 2024), which is well-known to be the asymptotic covariance matrix of the MLE. It follows that under correct specification of the polychoric model, the generalized ML estimator is indeed asymptotically first and second order equivalent to the consistent and efficient MLE of Olsson, (1979).
A consistent estimator of the unobserved asymptotic covariance matrix $\bm{\Sigma}\left({\bm{\theta}_{0}}\right)$ can be constructed as follows. Replace all population class probabilities $f_{\varepsilon}\left(x,y\right)$ by their corresponding empirical counterparts $\widehat{f}_{N}(x,y)$ in matrices $\bm{W}\left(\bm{\theta}\right),\bm{M}\left(\bm{\theta}\right)$, and $\bm{\Omega}$. Then exploit the plug-in principle and evaluate $\bm{U}\left(\bm{\theta}\right)$ and $\bm{M}\left(\bm{\theta}\right)$ at the point estimate $\widehat{\bm{\theta}}_{N}.$ Denote the ensuing plug-in estimator by $\bm{\Sigma}\left({\widehat{\bm{\theta}}_{N}}\right)$, which is consistent for $\bm{\Sigma}\left({\bm{\theta}_{0}}\right)$ by Theorem 1 and the continuous mapping theorem.
For the choice $c=0$, note that the condition $\frac{f_{\varepsilon}\left(x,y\right)}{p_{xy}(\bm{\theta}_{0})}-1\neq c$ in Theorem 1 rules out the zero-misspecification case ($\varepsilon=0$). Indeed, Welz, (2024) shows that for $c=0$, the estimator is not asymptotically Gaussian if the model is correctly specified. We therefore recommend to choose $c>0$ (see Section 5.3 for a discussion).

### v4

@@ 8.2 Simulation study
To simulate ordinal variables that were generated by a nonnormal latent distribution $G$, we employ the VITA simulation method of Grønneberg & Foldnes, (2017). For a pre-specified value of the population correlation $\rho_{G}$, the VITA method models the latent random vector $(\xi,\eta)$ such that the individual variables $\xi$ and $\eta$ both possess standard normal marginal distributions with population correlation set equal to $\rho_{G}=\mathrm{Cor}_{G}\left[\xi,\ \eta\right]$, but are not jointly normally distributed. Instead, their joint distribution $G$ is equal to a pre-specified nonnormal copula distribution, such as the Clayton or Gumbel copula. Grønneberg & Foldnes, (2017) show that discretizing such VITA latent variables yields ordinal observations that could not have been generated by a standard bivariate normal distribution, thereby ensuring proper violation of the latent normality assumption.
@@
such that both resulting ordinal variables have five response options each. We generate $N=1,000$ ordinal responses according to this data generation process and compute across 5,000 repetitions the same estimators and performance measures as in the simulations in Section 6.
@@
Figure 9 visualizes the bias of the robust estimator and the polychoric MLE under both Clayton copulas across the repetitions. [^15] In 27 of the 5,000 repetitions for the Clayton copula, the robust estimator experienced numerical instability and subsequently did not converge to a solution. The polychoric MLE failed to converge one time. Consequently, these unconverged estimates are omitted from Figure 9. We explain this numerical issue in detail in Appendix B. For correlation 0.9, the polychoric MLE exhibits a noteworthy bias, whereas the robust estimator remains accurate, albeit with a larger estimation variance as compared to partial misspecification (cf. Section 6). Conversely, for the weaker correlation 0.3, both estimators are fairly accurate with average biases of about 0.025-0.025- 0.025. Table 5 contains additional performance measures regarding inference. For $\rho_{G}=0.3$, the average standard error estimate of the robust method is accurate, but for $\rho_{G}=0.9$, it notably overestimates. We therefore also computed the median of the standard error estimates in the latter case: at 0.059, it still overestimates the approximate true standard error of 0.037, but far less so than the average. It turns out that there is a small number of simulated datasets with a large majority of empty cells in the contingency table, resulting in numerical instability of the standard error estimates and inflating their average. As discussed in Section 4.3, distributional misspecification is not covered by our partial misspecification framework, so it is not surprising that standard errors derived under partial misspecification are not always valid. Bootstrap inference may therefore be an attractive alternative. Nevertheless, unlike the polychoric MLE with zero coverage for $\rho_{G}=0.9$, the robust estimator maintains high coverage of over 90%.
Figure 9: Boxplot visualization of the bias of the robust estimator and the polychoric MLE, $\widehat{\rho}_{N}-\rho_{G}$, under distributional misspecification via a Clayton copula with correlation $\rho_{G}=0.9$ (left) and $\rho_{G}=0.3$ (right), across 5,000 repetitions. Diamonds represent the respective average bias. The tuning constant of the robust estimator is set to $c=0.6$.
Point estimate Standard error Confidence interval
Correlation Estimator $\widehat{\rho}_{N}$ Bias SE $\widehat{\text{SE}}$ Bias Coverage Length
$\rho_{G}=0.9$ Robust 0.893 0.007-0.007- 0.007 0.037 0.264 0.226 0.930 1.034
ML 0.799 0.101-0.101- 0.101 0.018 0.019 0.001 0.000 0.075
$\rho_{G}=0.3$ Robust 0.278 0.022-0.022- 0.022 0.039 0.039 0.000 0.916 0.154
ML 0.271 0.029-0.029- 0.029 0.037 0.037 0.000 0.885 0.146
Table 5: Results for the robust estimator with $c=0.6$ and the polychoric MLE across 5,000 simulated datasets under distributional misspecification via a Clayton copula with true population correlation $\rho_{G}\in\{0.9,0.3\}$. See Table 1 for explanatory notes on the performance measures.
@@
These results suggest that at least for point estimation, the Clayton copula with correlation 0.9 might be reasonably approximable by a contaminated normal distribution where the normal distribution covers the center of the probability mass and some contamination distribution covers the tails (cf. Section 8.1). Indeed, Figure 8 indicates that the normal distribution and Clayton copula at correlation 0.9 seem to behave similarly in the center, but deviate from one another towards the tails. On the other hand, at correlation 0.3, the two densities do not appear to be drastically different from one another, which may explain why both the polychoric MLE and the robust estimator work reasonably well for such distributional misspecification. If the latent distribution is not too different from a normal distribution, then the polychoric model may offer a satisfactory fit despite being technically misspecified.
@@
Overall, this simulation study demonstrates that in some cases of distributional misspecification, robustness can be gained with our estimator, compared to the polychoric MLE. However, it also demonstrates that there are cases of distributional misspecification for which the polychoric MLE still works quite well such that the robust estimator offers little gain. Nevertheless, the fact that in some situations robustness can be gained under distributional misspecification represents an overall gain in robustness.
@@ 9 Discussion and conclusion
We consider a situation where the polychoric correlation model is potentially misspecified for a subset of responses in a sample, that is, a set of uninformative observations not generated by a latent standard normal distribution. This model misspecification framework, called partial misspecification here, stems from the robust statistics literature, and a relevant special case is that of careless respondents in questionnaire studies. We demonstrate that maximum likelihood estimation is highly susceptible to the presence of such uninformative responses, resulting in possibly large estimation biases and low coverage of confidence intervals.
As a remedy, we propose an estimator based on the $C$-estimation framework of Welz, (2024) that is designed to be robust to partial model misspecification. Our estimator generalizes maximum likelihood estimation, does not make any assumption on the magnitude or type of potential misspecification, comes at no additional computational cost, and possesses attractive statistical guarantees such as asymptotic normality. It furthermore allows to pinpoint the sources of potential model misspecification through the notion of Pearson residuals (PRs). Each possible response option is assigned a PR, where values substantially larger than the ideal value 0 imply that the response in question cannot be fitted well by the polychoric correlation model. In addition, the methodology proposed in this paper is implemented in the free open source package robcat (Welz et al.,, 2025) for the statistical programming environment R and is publicly available at https://CRAN.R-project.org/package=robcat.
@@
Although not covered by our partial misspecification framework, we also discuss how and when our estimator can offer a robustness gain (compared to the polychoric maximum likelihood estimator) when the polychoric model is misspecified for all observations, which has been a subject of interest in recent literature. In essence, there can be a robustness gain if the latent nonnormal distribution that generated the data can be reasonably well approximated by a contaminated normal distribution where the normal distribution reflects the central part and some unspecified contamination distribution reflects the tails.
@@
We verify the enhanced robustness and theoretical properties of our robust estimator in simulation studies. Furthermore, we demonstrate the estimator’s practical usefulness in an empirical application on a Big Five administration, where we find compelling evidence for the presence of careless respondents as a source of partial model misspecification.
@@
However, our estimator depends on a user-specified choice of a tuning constant $c$, which governs a tradeoff between robustness and efficiency in case of misspecification. While simulation experiments suggest that the choice $c=0.6$ provides a good tradeoff and estimates do not change considerably for a broad range of finite choices of $c$, a detailed investigation on this tuning constant needs to be carried out in future work. As an alternative, one could consider other discrepancy functions that do not depend on tuning constants, like the ones discussed in Welz, (2024).
A practical consideration of polychoric correlation is computing time. Our estimator simultaneously estimates all model parameters (i.e., correlation coefficient and thresholds) for robustness reasons, hence it is computationally more intensive than the non-robust two-stage approach of Olsson, (1979). To alleviate the computational burden, our implementation in package robcat is written in fast and efficient C++ code. Furthermore, its default behavior first tries a fast unconstrained numerical optimization routine, which in our experience almost always suffices and executes in about half a second for five-point rating variables on a regular laptop. For estimating polychoric correlation matrices, package robcat supports parallel computing to keep computation time low. Since our method generalizes maximum likelihood and has the same time complexity, these functionalities also provide a fast implementation of the maximum likelihood estimator.
@@
The methodology proposed in this paper suggests a number of extensions. For instance, one could use a robustly estimated polychoric correlation matrix in structural equation models to robustify such models against misspecification. Similar robustification could by achieved in, for instance but not limited to, principal component analyses, multidimensional scaling, or clustering. In addition, the theory of Welz, (2024) may allow to pinpoint possible sources of model misspecification on the individual response level. That is, it may enable the derivation of statistically sound cutoff values for Pearson residuals in order to detect whether a given response can be fitted well by the polychoric correlation model. We leave these avenues to further research.
@@ Appendix A Asymptotic properties of the robust estimator
This appendix section presents the limit theory of the proposed robust estimator $\widehat{\bm{\theta}}_{N}$ in (10). Throughout this section, we assume that the number of response categories, $K_{X}$ and $K_{Y}$, are fixed and known, and that the sample $\{(X_{i},Y_{i})\}_{i=1}^{N}$ used to compute an estimate $\widehat{\bm{\theta}}_{N}$ has been generated by the process in (1) where the latent $\{(\xi_{i},\eta_{i})\}_{i=1}^{N}$ are draws from distribution $G_{\varepsilon}$ in (7) with unobserved contamination fraction $\varepsilon\in[0,0.5)$ and unknown contamination distribution $H$.
@@ A.1 Main theorem
We start by introducing additional notation. Denote the likelihood score function, i.e., the $d$-dimensional gradient vector of $\log\left(p_{xy}(\bm{\theta})\right)$ for response $(x,y)\in\mathcal{X}\times\mathcal{Y}$ at parameter $\bm{\theta}\in\bm{\Theta}$, by
$\bm{s}_{xy}(\bm{\theta})=\frac{\partial}{\partial\bm{\theta}}\log\left(p_{xy}(\bm{\theta})\right)=\frac{1}{p_{xy}(\bm{\theta})}\left(\frac{\partial}{\partial\bm{\theta}}\ p_{xy}(\bm{\theta})\right).$
Further, given the tuning constant $c\in[0,\infty]$, define the $d\times d$ matrix
$\bm{Q}_{xy}(\bm{\theta})=\Bigg{(}\frac{f_{\varepsilon}\left(x,y\right)}{p_{xy}(\bm{\theta})}\mathds{1}\left\{\frac{f_{\varepsilon}\left(x,y\right)}{p_{xy}(\bm{\theta})}-1\in[-1,c]\right\}+(c+1)\mathds{1}\left\{\frac{f_{\varepsilon}\left(x,y\right)}{p_{xy}(\bm{\theta})}-1\ >c\right\}\Bigg{)}\frac{\partial^{2}}{\partial{\bm{\theta}}\partial{\bm{\theta}}^{\top}}p_{xy}(\bm{\theta})$
@@
For $c\in[0,\infty]$ the tuning constant in the discrepancy function, assume that $\frac{f_{\varepsilon}\left(x,y\right)}{p_{xy}(\bm{\theta}_{0})}-1\neq c$ for all $(x,y)\in\mathcal{X}\times\mathcal{Y}$. Then, under certain regularity conditions that do not restrict the degree or type of possible misspecification of the polychoric model beyond $\varepsilon\in[0,0.5)$, when $N\to\infty$ it holds true that
@@
$\bm{U}\left(\bm{\theta}\right)$ $=\mathbb{V}\mathrm{ar}_{f_{\varepsilon}}\left[\bm{w}_{XY}(\bm{\theta})\right]=\bm{W}\left(\bm{\theta}\right)\bm{\Omega}\bm{W}\left(\bm{\theta}\right)^{\top}\qquad\text{and}$
$\bm{M}\left(\bm{\theta}\right)$ $=\frac{\partial^{2}}{\partial{\bm{\theta}}\partial{\bm{\theta}}^{\top}}L\big{(}\bm{\theta},f_{\varepsilon}\big{)}$
$=\sum_{x\in\mathcal{X}}\sum_{y\in\mathcal{Y}}\Bigg{[}\mathds{1}\left\{\frac{f_{\varepsilon}\left(x,y\right)}{p_{xy}(\bm{\theta})}-1\in[-1,c]\right\}f_{\varepsilon}\left(x,y\right)\bm{s}_{xy}(\bm{\theta})\bm{s}_{xy}(\bm{\theta})^{\top}-\bm{Q}_{xy}(\bm{\theta})\Bigg{]}$
@@
It can be shown that in the absence of misspecification (such that $\bm{\theta}_{0}=\bm{\theta}_{*}$), the asymptotic covariance matrix $\bm{\Sigma}\left({\bm{\theta}_{*}}\right)$ is equal to the inverted Fisher information matrix of the polychoric model (Lemma B.5 in Welz,, 2024), which is well-known to be the asymptotic covariance matrix of the MLE. It follows that under correct specification of the polychoric model, the proposed robust estimator is indeed asymptotically first and second order equivalent to the consistent and efficient MLE of Olsson, (1979).
It can furthermore be shown that Theorem 1 generalizes well-known results on the behavior of ML estimation in misspecified models. Indeed, for the MLE ($c=+\infty$ in the discrepancy function), the sandwich-type asymptotic covariance matrix $\bm{\Sigma}\left({\bm{\theta}_{0}}\right)$ reduces to that of White, (1982, Theorem 3.2) and Huber, (1967, Corollary on p. 231). However, our asymptotic results are more general because they encompass a broad class of estimators, including and beyond ML.
A consistent estimator of the unobserved asymptotic covariance matrix $\bm{\Sigma}\left({\bm{\theta}_{0}}\right)$ can be constructed as follows. Replace all population class probabilities $f_{\varepsilon}\left(x,y\right)$ by their corresponding empirical counterparts $\widehat{f}_{N}(x,y)$ in matrices $\bm{W}\left(\bm{\theta}\right),\bm{M}\left(\bm{\theta}\right)$, and $\bm{\Omega}$, and denote the resulting observable matrices by $\widehat{\bm{W}}_{N}\left(\bm{\theta}\right),\widehat{\bm{M}}_{N}\left(\bm{\theta}\right)$, and $\widehat{\bm{\Omega}}_{N}$, respectively. Each of these observable matrices is a (pointwise) consistent estimator of its corresponding population counterpart by the continuous mapping theorem. Then apply the formulas in Theorem 1 to construct
$\widehat{\bm{\Sigma}}_{N}\left(\bm{\theta}\right)=\widehat{\bm{M}}_{N}\left(\bm{\theta}\right)^{-1}\widehat{\bm{U}}_{N}\left(\bm{\theta}\right)\widehat{\bm{M}}_{N}\left(\bm{\theta}\right)^{-1},$
where $\widehat{\bm{U}}_{N}\left(\bm{\theta}\right)=\widehat{\bm{W}}_{N}\left(\bm{\theta}\right)\widehat{\bm{\Omega}}_{N}\widehat{\bm{W}}_{N}\left(\bm{\theta}\right)^{\top}$. The observable matrix $\widehat{\bm{\Sigma}}_{N}\left(\bm{\theta}\right)$ is a pointwise consistent estimator of $\bm{\Sigma}\left({\bm{\theta}}\right)$ for a given parameter value $\bm{\theta}\in\bm{\Theta}$. It now follows from the fact that $\widehat{\bm{\theta}}_{N}\stackrel{{\scriptstyle\mathbb{P}}}{{\longrightarrow}}\bm{\theta}_{0}$ (Theorem 1) and the continuous mapping theorem that $\widehat{\bm{\Sigma}}_{N}\left(\widehat{\bm{\theta}}_{N}\right)$ is consistent for the population asymptotic covariance matrix $\bm{\Sigma}\left({\bm{\theta}_{0}}\right)$.
For the choice $c=0$, note that the condition $\frac{f_{\varepsilon}\left(x,y\right)}{p_{xy}(\bm{\theta}_{0})}-1\neq c$ in Theorem 1 rules out the zero-misspecification case ($\varepsilon=0$). In fact, in this case, the condition is not met for any $(x,y)\in\mathcal{X}\times\mathcal{Y}$. Indeed, Welz, (2024) shows that for $c=0$, the estimator is not asymptotically Gaussian if the model is correctly specified. We therefore recommend to choose $c>0$ (see Section 5.3 for a discussion). The reason why we require $\frac{f_{\varepsilon}\left(x,y\right)}{p_{xy}(\bm{\theta}_{0})}-1\neq c$ for all $(x,y)\in\mathcal{X}\times\mathcal{Y}$ is because the value $c$ is the threshold at which the discrepancy function (11) transitions from superlinear growth to linear growth. The discrepancy function is not twice differentiable at this point, which causes a particular Taylor expansion in the theorem’s proof to fail to exist. This Taylor expansion plays a crucial role in establishing asymptotic normality. If it cannot be performed for at least one response $(x,y)\in\mathcal{X}\times\mathcal{Y}$, the estimator is not asymptotically normal. We refer to Welz, (2024) for a detailed discussion.